OpenStudy (anonymous):

m(x)={8x x<2 x^2-1 x> or equal to 2 lim m(x)= x arrow>2-

OpenStudy (tkhunny):

Is the limit the same from the right and from the left?

OpenStudy (anonymous):

yes but we are only looking at the negative side 2^-

OpenStudy (tkhunny):

Oh, then this is continuous? Or isn't it?

OpenStudy (anonymous):

I think so then is the answer positive infinty

OpenStudy (tkhunny):

?? 8(0) = 0 8(1) = 8 8(3/2) = 12 8(7/4) = 14 8(15/8) = 15 8(31/16) = 15.5 8(63/32) = 15.75 8(127/64) = 15.875 Definitely not headed for "infinity".

OpenStudy (anonymous):

ok so then do youknow the answer?

OpenStudy (tkhunny):

You said it was continuous. Do we KNOW that or did you just assume it?

OpenStudy (tkhunny):

The left limit is 16 The right limit is 3 It's not continuous.